OpenStudy (anonymous):

Helppppp! draw a graph or just a line for this equation please? i cant figure out which way it goes. -6/7x -1/2y=3/7

OpenStudy (anonymous):



OpenStudy (anonymous):

Looks like this:

OpenStudy (anonymous):

In slope-intercept-form, that equation is: y = (-12/7)x - 6/7

OpenStudy (anonymous):

You can graph that by getting the x and y intercepts and then just draw a straight line through the points.

OpenStudy (anonymous):

THANK YOU BOTH! Can either of youu help me with this one too? –0.2x – 0.6y = 0.12

OpenStudy (anonymous):

You can get the y-intercept by seting "x" = to 0: y = (-12/7)(0) - 6/7 y = -6/7 Sure, fire away.

OpenStudy (anonymous):

You can get the x-intercept by setting "y" = 0: 0 = (-12/7)x - 6/7 x = -1/2

OpenStudy (anonymous):

btw, good luck with all of your studies and thx for the recognition!

OpenStudy (anonymous):

so what would that look like? and thank you for the help!

OpenStudy (anonymous):

The graph?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In my post under your drawing I have an attachment.

OpenStudy (anonymous):

No, I mean for the second equation. Should I make it a seperate question so you can get another meal? Its this one. –0.2x – 0.6y = 0.12

OpenStudy (anonymous):

metal*

OpenStudy (anonymous):

medal* omg Hhaha

OpenStudy (anonymous):

lol sorry, didn't see that one. Ok, I'll get to work. Gotta get the eyes working first! :-) That's okay, I'll answer it here. Medals are nice, but I'm here to help. Medals are way secondary.

OpenStudy (anonymous):

Awhhh thanks!

OpenStudy (anonymous):

So, here's a FRESH graph. Hot off of the computer:

OpenStudy (anonymous):

That equation I re-wrote to: y = -x/5 - 0.2

OpenStudy (anonymous):

x and y intercepts are found the same way as described for the other one.

OpenStudy (anonymous):

Hope you come back and see your spanking-new graph! It was nice working with you. I had to go get a cake out of the oven, so hopefully I'll be back later. If my cake didn't burn. After I scrape burnt cake off of the walls.

OpenStudy (anonymous):

That is amazing! Thank you sooooooo much!

OpenStudy (anonymous):

and lol, good luck with your cake! Merry Christmas! or Happy Hanukah! or Happy Kwanza! Whatever your little heart desires! haha

OpenStudy (anonymous):

You're quite welcome, @lyzzy7 and the cake turned out OK! Homemade banana cake from scratch, from me! A non-baker! And merry X-mas to you, too!